OpenStudy (erniellyd28):

recall that to angles are complementary. if the sum of their measures is 90 degrees. find the measures if one angle is twenty times the other angle.

OpenStudy (rivershinesdragon):

Let's draw it better, using the variable x: 

OpenStudy (rivershinesdragon):

20x + x = 90. So, solve for x is only step 1. What is x?

OpenStudy (erniellyd28):

70?

OpenStudy (rivershinesdragon):

Wait... X is not 70, do you mean the angle 20x?

OpenStudy (erniellyd28):

no im very sorry i was confused i thought the answer is 70

OpenStudy (rivershinesdragon):

It's fine; I'm here to help. Just gimme a sec...

OpenStudy (rivershinesdragon):

\[20x +x =90\] \[\frac{ 21x }{ 21 } = \frac{ 90 }{ 21 }\]

OpenStudy (rivershinesdragon):

\[x \approx 4\]

OpenStudy (rivershinesdragon):

So, step 2: substitute x in each angle. Can you solve for 20x given the information?

OpenStudy (erniellyd28):

no

OpenStudy (rivershinesdragon):

Well, here's a hint: 

OpenStudy (rivershinesdragon):

What's 20*4?

OpenStudy (erniellyd28):

80

OpenStudy (sshayer):

\[x=\frac{ 90 }{ 21 }=\frac{ 30 }{ 7 }\] 20x=?

OpenStudy (erniellyd28):

theres no number that goes into 30 and 7 equally

OpenStudy (rivershinesdragon):

That's why you leave it as 30/7

OpenStudy (rivershinesdragon):

So, x = 30/7, which is very close to 4. \[30/7 * 21 = 90\]

OpenStudy (rivershinesdragon):

4 and 30/7 are acceptable answers. Good job

OpenStudy (sshayer):

\[\frac{ 30 }{ 7 }=4\frac{ 2 }{ 7 }\] it is a real number.

OpenStudy (sshayer):

\[20x=20*\frac{ 30 }{ 7 }=?\]

OpenStudy (erniellyd28):

but now im lost as to which number i should use as my larger angle and which i should use as my smaller angle

OpenStudy (sshayer):

it is understood 20x>x,x being positive.

OpenStudy (sshayer):

second angle=20x=?

OpenStudy (sshayer):

\[20x=20*\frac{ 30 }{ 7 }=\frac{ 600 }{ 7 }=85\frac{ 5 }{ 7 }\]

OpenStudy (erniellyd28):

now im even more confused lol

OpenStudy (erniellyd28):

im very sorry

OpenStudy (sshayer):

sum of these angles=90 you can check.

OpenStudy (sshayer):

\[\frac{ 30 }{ 7 }+\frac{ 600 }{ 7 }=\frac{ 30+600 }{ 7 }=\frac{ 630 }{ 7 }=90\]

OpenStudy (erniellyd28):

im lost but ill figure it out thank you both very much for your help :-)

OpenStudy (sshayer):

can you tell me where is the problem?

OpenStudy (sshayer):

x is not 70 it is like we say 20 apples+1 apple=21 apples similarly 20x+1x=21x